John deposits 10000 in a college trust fund in which i=6.4% compounded quarterly. How much $ would be in the account after 18 yrs.




#2 In an account which pays 6.5% compounded continuosly, how much $ would need to be deposited in order to rach 50000 in 15 yrs?

Answer by stanbon(75887)

John deposits 10000 in a college trust fund in which i=6.4% compounded quarterly. How much $ would be in the account after 18 yrs.
-----------
A(t) = P(1+(r/n))^(nt)
A(18) = 10,000(1+(0.064/4))^(4*18)
A(18) = 10,000(1+0.016)^72
A(18) = 10,000*3.136
A(18) = $31,360
===========================
#2 In an account which pays 6.5% compounded continuosly, how much $ would need to be deposited in order to rach 50000 in 15 yrs?
---
A(t) = Pe^(rt)
50,000 = Pe^(0.065*15)
50,000 = P(2.651)
P = $18,859.62
=====================
Cheers,
Stan H.